OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y is less than or equal to x-5 y is less than or equal to -x-4 a) (1,10) b) (-1, 10) c) (10, 1) d) (1, -10)

Parth (parthkohli):

Where do you have a problem? I'd be elated to help you.

OpenStudy (anonymous):

What must I do to solve the problem?

Parth (parthkohli):

You may try these: 1) \( \color{Black}{\Rightarrow 10 \le 1 - 5? }\) \( \color{Black}{\Rightarrow 10 \le -1 - 4?}\) 2) \( \color{Black}{\Rightarrow y \le -1 - 5?}\) \( \color{Black}{\Rightarrow y \le 1 - 5? }\)

Parth (parthkohli):

Remember: A solution set is just a pair of numbers that \(\textbf{satisfy}\) an inequality.

Parth (parthkohli):

Also, a solution set is in the form \((x,y)\).

OpenStudy (anonymous):

But what is the first step that I must take to solve this problem?